how many moles of hydrochloric acid are dissolved in 500 mL of a 6.3 solution? Show all work.


1
Expert's answer
2022-04-08T07:42:05-0400

C(sol) = n(HCl)/V(sol)

n(HCl) = V(sol) × C(sol) = 0.500 L × 6.3 mol/L = 3.15 mol